Let x>0, and y be a positive function of x such that x^y=y^-x.

Find y′



1
Expert's answer
2021-10-21T14:45:24-0400

"x^{y}=y^{(-x)}"

Taking logerithm with respect to base e on both sides

"\ln{x^{y}}=\ln{y^{(-x)}}"

=> "y\ln{x}=-x\ln{y}"

Differentiating with respect to x

"\frac{d(y\ln{x})}{dx}= - \frac{d(x\ln{y})}{dx}"

=> "y\frac{d}{dx} (\ln{x})+\ln{x} \frac{dy}{dx}=-x\frac{d}{dx}(\ln{y})-\ln{y} \frac{d}{dx}{(x)}"

=> "y.\frac{1}{x} + ln{x} \frac{dy}{dx}=-x.\frac{1}{y}.\frac{dy}{dx}-\ln{y}"

=> "\ln{x} \frac{dy}{dx}+x.\frac{1}{y}.\frac{dy}{dx}" = "-\frac{y}{x} - \ln{y}"

=> "x(y\ln{x}+x)\frac{dy}{dx}= -y(y+x\ln{y})"

=>

=> "\frac{dy}{dx}=" "-\frac{ y(y+x\ln{y})}{x(y\ln{x}+x)}"

So , y' = "-\frac{ y(y+x\ln{y})}{x(y\ln{x}+x)}"